If the volume of sales is $7,000,000 and sales at the break-even point amount to $4,800,000, the margin of safety is 45.8%.

True

False
Answer #1

margin of safety=Total sales-Breakeven sales

=(7,000,000-4,800,000)=2,200,000

margin of safety %=margin of safety/Total sales

=2,200,000/7,000,000

=31.4%(Approx)

Hence the statement is False.